# PooDLe[Uncaptioned image]: Pooled and dense self-supervised learning from naturalistic videos

### v1

@@ 4.4 Spatial and temporal cropping in self-supervised video learning
In this section, we study the effect of image cropping parameters used during data augmentation. As there is no longer a 1:1 image-to-concept relationship like in iconic data, how much of a scene is visible and assigned to each representation can greatly affect learning. Specifically, we perform 5 experiments varying: (1) global crop area, (2) subcrop area, and (3) global crop area and resolution to preserve pixel density (4) number of subcrops (5) temporal stride $\Delta t$. Crop area refers to the fraction of an image taken by the crop, and pixel density is the ratio of source frame pixels to model input frame pixels. Figure 6 visually depicts how crops transition from global to pseudo-iconic with decreasing crop area. Training recipe follows the ablations in section 4.3 and should be compared to row 7 in Table 4.2.
@@
Figure 9: mIoU when varying $\Delta t$.
Varying global crop area when training PooDLes.
We train 4 PooDLes on non-overlapping crop ranges $(a_{\text{min}},a_{\text{max}})=(0.05,0.25),(0.25,0.50),(0.50,0.75),(0.75,1.0)$, and 2 different resolutions $256\times 512$ and $384\times 768$. Results are shown in Figure 5. We find that directly increasing crop area, thereby increasing the content seen by each vector in $\mathcal{L}_{\text{dense}}$, does not directly improve performance. Rather, there is a sweet spot for each resolution, $[0.05,0.25]$ for $256\times 512$ and $[0.25,0.5]$ for $384\times 768$ with the larger resolution performing slightly better (34.0% v. 34.4%). While larger global crop areas improve PooDLe representations, they must be paired with a larger input resolution, otherwise performance degrades. We believe this is due to loss of information from downsampling.
@@
Next, we study how subcrop area affects our learned representations. Again, we train 4 PooDLes on subcrop ranges, using a fixed global crop range [0.125, 0.25] at resolution $256\times 512$. Results are shown in Figure 5. We show the mIoU for all classes, as well as the large and small class subgroupings. We observe that larger subcrop areas result in poorer performance, which we hypothesize is due to how larger subcrops are more likely to contain multiple subjects, producing false invariances. In addition, when crops are too large, subcrops may not serve well as pseudo-iconic views.
Varying input resolution with fixed pixel density.
With the above observation, we train another set of models while controlling the pixel density. The frame crop area is set to $[0.125,0.25]$ and the base crop size, $256\times 512$ making the average pixel density $\rho_{\text{pixel}}=1.32$. Maintaining this pixel density, we extrapolate 3 additional scales with input resolution $H\times W=(192\times 384),(320\times 640),(384\times 768)$ and crop range $(0.07,0.14),(0.195,0.39),(0.281,0.563)$. We pretrain a PooDLe as well as dense-only, pooled-only, and FlowE models for each scale on BDD100K and evaluate using linear readout. As subcrops are taken from global crops, we adjust the local crop area following Figure 5 for $320\times 640$ and $384\times 768$ to $(0.032,0.192)$ and $(0.022,0.133)$.
The results in Figure 5 confirm our hypotheses. PooDLe trained with both objectives is more robust to different input scopes given a fixed pixel density whereas a dense-only model requires high resolution, and pooled-only and FlowE are less competitive on all resolutions.
Varying number of subcrops
We also study how varying the number of subcrops affects performance. We train 4 PooDLes using $K=0,3,6,9$ subcrops on BDD100K and evaluate using linear readout, with results shown in Figure 7. Using 3 subcrops gives an initial large performance jump and using additional subcrops provides more modest gains. We decide to use $K=6$ as our default option to balance between performance and computational efficiency.
@@
To study the effect of varying $\Delta t$, we train PooDLe with values $\Delta t=0,15,30,45$ on BDD100K and once again evaluate using linear readout (Figure 9). Performance degrades when temporal stride is either too small or too large. When it is small, there is limited variance in object appearance, making the training objective too easy. On the other hand, as temporal stride grows, correspondence between frames decreases and optical flow may become unreliable. Note that for $\Delta t=0$, we jitter the initial large crop by up to $10\%$ of the image size. We also visualize frame sequences from 3 different videos, Figure 9 showing the high variability of motion in BDD100K.
@@ 4.5 Subcrops as pseudo-iconic training images
We find that the addition of a pooled loss on pseudo-iconic subcrops improves PooDLe representations, particularly on smaller objects. To better understand this effect, we analyze how subcrops may increase the prevalence of small objects when they are assumed to be pseudo-iconic. First, we simulate a circle as our foreground object and compute the probability that a subcrop “hits” the object as a pseudo-iconic view, as illustrated in Figure 10(a). A subcrop counts as a hit if at least 5% of the subcrop is occupied by the object. We argue that this is a reasonable assumption because background classes typically contain little visual variation and thus, have a relatively modest effect on the learned pooled representations. We also empirically simulate subcrops on the BDD100K semantic segmentation training dataset, and compute subcrop hit and pixel probabilities for foreground objects of varying size. As shown in Figure 10(b), the relative difference between the subcrop probability and pixel probability is greater for smaller objects. This illustrates how subcrops can boost prevalence of small foreground objects in the pooled objective in comparison to the pixel-level dense SSL objective. Appendix C contains additional details about our setup and analyses.
5 Conclusion
(a) Toy simulation to compute probability of subcrop acting as a pseudo-iconic view, i.e. hit, of circular object. Subcrop hits are green and non-hits are red.
@@ 5 Conclusion
Self-supervised learning on naturalistic videos present many unsolved challenges, especially due to the presence of high-resolution multi-object crowded scenes with severe spatial region imbalance. Iconic methods rely on single-subject images and dense methods struggle with scale-imbalance of objects. We propose PooDLe that combines pooled region-invariance learning and dense flow-equivariance learning objectives in a unified framework. PooDLe achieves state-of-the-art performance on downstream semantic evaluations compared to prior methods pretrained on the same video datasets, particularly on recognizing small objects. Our study on the effects of crop area, resolution density, and temporal stride also offers key insights on the design choices for video self-supervised learning.
Limitations and future work.
Our work has several limitations and areas for future work. First, PooDLe relies on an external flow prediction network. Although we demonstrate that an unsupervised flow predictor also works well, it would be more elegant to incorporate motion prediction in a joint framework like in MC-JEPA [5]. Additionally, while PooDLe achieves state-of-the-art performance, we find that it does not currently work well with ViTs and this limitation warrants further investigations. In the future, we would also like to explore ways to zoom-in and attend to interesting regions and regions that contain rare objects rather than relying on random subcrops, e.g., using priors such as motion [13, 28] to identify foreground objects.
Acknowledgment
We would like to thank Adrien Bardes and Jean Ponce for helpful discussion, and the Microsoft Accelerating Foundation Models Research program for providing cloud compute credits for our experiments. The project was also supported by the NYU High Performance Computing resources, services, and staff expertise.
@@ Appendix A Implementation details
As discussed in the pretraining details, we use a Resnet-50 [23] as our backbone architecture. The projector model is a non-linear, 2-layer MLP (linear for pooled, $1\times 1$ convolutions for dense) that has a $4096$ hidden dimension and projects down to $256$ dimensions. The predictor is the same network with $256-4096-256$ channels. We follow BYOL [20] with a momentum starting at $0.996$ and increases to $1$ throughout training.
@@
Flow is predicted using a supervised off-the-shelf RAFT model or an unsupervised UFlow [30] model that we train ourselves. For unsupervised training, we exactly follow UFlow and train on the KITTI [18] dataset before finetuning on BDD100k [50] for 100000 steps on daytime-only videos. The training and inference resolutions were set to $256\times 512$ to better match the inference setting. KITTI used adjacent frames (10Hz video) while BDD frames were sampled with temporal stride of 10 (30Hz video).
@@
$K=6$ paired local crops are sampled using the methods described. Cropping is performed using RandomResizedCrop with an output resolution of $192\times 192$. Jitter is 10% of the input image size and standard aspect ratio range of $[3/4,4/3]$ is used.
@@
We sum our 2 loss functions directly and give them equal weight. The loss computation and warping function were applied to representations after reversing the affine transform and resizing to the input image resolution, this is to take full advantage of high resolution flow like in FlowE [49]. We also use flow-based occlusion to prevent misaligning occluded regions without correspondence. We use the same occlusion formulation as DDFlow [32] and parameters $\alpha_{1}=0.1,\alpha_{2}=0.5$. We also mask out regions that are not visible after affine transformations for $\mathcal{L}_{\text{dense}}$.
@@
AdamW is used as the optimizer and a weight decay value of $0.01$. A learning rate of $45e-$ 4 is used with batch size of 32 GPUs and $4$ image pairs per GPU for a batch size total of $128$. Cosine learning rate decay is used with a schedule for 300 epochs, despite early termination due to compute limitations. LR warmup is used for 2 training epochs. Full float32 precision is used during training.
@@
For all BDD and Cityscapes semantic segmentation and object detection readout tasks, we follow the setup described in FlowE [49] for ResNet-based methods. For ViT-based methods, we adopt those settings, but use AdamW for the optimizer with learning rate of $53e-$ 5 and weight decay of $0.05$, and a crop size of $512\times 512$ rather than the normal $512\times 1024$ to accommodate the square aspect ratio used in ViT-pretraining, following the semantic segmentation linear readout setup described in iBOT [54]. In addition, ViT-based methods require sliding window inference in order to achieve performance that is competitive with convolution-based methods.
For ADE20K linear readout, we simply use the respective BDD linear readout settings for ResNet and ViT methods. For ADE20K UperNet finetuning, we follow the procedure described in iBOT [54].
Abaltion data sampling.
For all ablation experiments, we employ repeated sampling like in MAE-st [17] which samples $R$ frames each time a video is encountered for faster data loading. Therefore, each pass through every video in the dataset counts as $R$ epochs.
@@ Appendix B Compute resources
Ablation-sized experiments were run 2 or 4 H100/A100 GPUs for a total of 40 epochs taking 20-40h depending on the configuration.
@@ Appendix C Subcrop analysis
Appendix D Additional visualizations
We provide additional visualizations of results on our evaluated benchmarks: BDD100K [50] semenatic segmentation (Figure 12), object detection (Figure 13) and ADE20K semantic segmentation (Figure 14). Once again, we note that PooDLe produces segmentation maps with clearer boundaries while also effectively capturing small objects.
@@ Appendix D Additional visualizations
Appendix E Additional evaluation results
Table 5: Additional BDD semantic segmentation (SemSeg) and object detection (Det) readout evaluations. All settings are conducted with a frozen backbone. BYOL results are taken from FlowE [49] and used DeepLab v3 [9] in-place of Upernet [47]. *Pretrained on BDD, initialized with supervised ImageNet weights.
@@ Appendix E Additional evaluation results
BYOL [20] R50 1000 IN1K 28.3 - 52.4 - 2.8 26.0 39.9 - 60.3 -
DenseCL [45] R50 200 IN1K 21.3 82.7 52.8 91.6 0.3 25.0 27.3 84.0 63.7 93.7
@@
DINO [7] ViT-S 800 IN1K 38.5 88.1 52.3 92.0 - - 47.1 90.3 63.6 94.0
iBOT [54] ViT-S 800 IN1K 44.4 89.6 54.2 92.2 - - 52.1 91.5 65.3 94.3
@@
We compare PooDLe against ImageNet-pretrained baselines in Table 5 and observe that PooDLe outperforms most baselines except iBOT and ImageNet supervised ViT. This result is encouraging, as pretraining on naturalsitic video is more challenging due to spatial and class imbalance, yet is also a more realistic setting that enables the use of broader sets of usable data. Furthermore, we note that pretraining on class-balanced data such as ImageNet particularly benefits mIoU, which weighs all classes equally despite some classes only appearing in a tiny proportion of pixels in evaluation. Finally, PooDLe pretrained on BDD with weights initialized from the ImageNet supervised checkpoint surpasses all ImageNet-pretrained baselines on linear semantic segmentation.
Appendix F Per-class evaluation results
Table 6: IoU per class on BDD semantic segmentation linear readout.
@@ Appendix F Per-class evaluation results
Appendix G Spatial token collapse in vision transformers
Figure 15: Comparison of L2 norm of last feature layer activations. The visualizations show both the raw L2 norm values as well as the L2 norm values clamped at the 80th percentile.
Initial attempts at using a ViT [16] backbone for PooDLe results in poor downstream performance. We hypothesize this is because of unwanted interaction between the model architecture and the FlowE dense objective. In our initial analysis, we have found that ViTs trained with only the FlowE dense objective results in background tokens receiving very high feature activations relative to other regions, as shown in Figure 15. This does not occur for our ResNet-based model as well as DINO-ViT which we believe is contributing to the performance drop. In the future, we would like to investigate this phenomenon further and formulate possible remedies to enable PooDLe to work with ViTs.
Appendix H Flow visualizations
Figure 16: Comparison of predicted optical flow from RAFT (supervised) and UFlow (unsupervised)
In Figure 16, we compare the predicted flow maps generated from RAFT [40], an off-the-shelf supervised model, and our own unsupervised UFlow [30] model. The frame pairs are randomly sampled with $\Delta t\in[15,30]$. We do note that self-supervised flow, particularly on BDD100K may exhibit noisy or splotchy results. This is possibly due to the inconsistent motion and large dark regions that do not offer sufficient photometric supervisory signal. This is in contrast to RAFT [40] which learns sharp edges like from supervised labels. Nevertheless, we find that this self-supervised flow is sufficient for training PooDLes.

### v2

@@ 4.4 Spatial and temporal cropping in self-supervised video learning
In this section, we study the effect of frame intervals and image cropping parameters used in data augmentation. Without a 1:1 image-to-concept relationship like in iconic data, the visible area of each frame can greatly affect representation learning. To study this, we perform 4 experiments varying: (1) subcrop area, (2) global crop area, (3) number of subcrops, (4) temporal stride $\Delta t$. Crop and subcrop area refer to the fraction of the frame taken during random-resized cropping. Figure 6 depicts how crops transition from global to pseudo-iconic with decreasing crop area. Training recipe follows the ablations in section 4.3 and results should be compared to row 7 in Table 4.2.
@@
Figure 9:
@@
First, we study how subcrop area affects our learned representations. We train 4 PooDLes using different subcrop ranges and a fixed global crop area of [0.125, 0.25] at resolution $256\times 512$. Results are shown in Table 5 for all classes, as well as small and large class subgroupings. We observe that larger subcrop areas result in worse performance, with a larger relative drop for smaller classes. This is likely because larger crops contain multiple, smaller objects which breaks the pseudo-iconic assumption and produces false invariances.
Varying global crop area
We select 3 different resolutions, and sample the crop area from a Gaussian truncated to $[0,1]$, with varying mean and $\sigma=0.1$. All three resolutions are trained with $\mu=0.125,0.275,0.425,0.725$ with the 2 smaller resolutions also trained at $\mu=0.05,0.20$ for higher granularity. The largest model ($512\times 1024$) produces the best results and peaks at $\mu=0.425$ while the other two peak at smaller crop areas. The $256\times 512$ model degrades more slowly in performance as crop area increases in comparison to the $128\times 256$ model.
Varying number of subcrops.
We also study how varying the number of subcrops affects performance. We train 4 PooDLes using $K=0,3,6,9$ subcrops on BDD100K and evaluate using linear readout, with results shown in Figure 7. Using 3 subcrops gives an initially large performance jump, and using additional subcrops provides more modest gains. We decide to use $K=6$ as our default option to balance between performance and computational efficiency.
@@
When it is small, there is limited variance in object appearance, diminishing the value of video data, and when it is too large, correspondence between frames decreases and optical flow becomes unreliable. Note that for $\Delta t=0$, we add jitter to the initial large crop by up to $10\%$ of the image size. Figure 9 shows frame sequences from 2 different videos, highlighting the high variability of motion in BDD100K.
@@ 4.5 Subcrops as pseudo-iconic training images
(a) Toy simulation of the probability of subcrop acting hitting a circular object. Subcrop hits are green and non-hits are red.
@@
To understand the impact of pseudo-iconic subcrops on PooDLe’s performance, particularly on small objects, we analyze their effect on object prevalence in the pooled objective. Using a simulated circular object (Figure 10(a)), we calculate the probability of a subcrop capturing it as a pseudo-iconic view, i.e. subcrop “hit”, and compare this to the probability of a pixel landing on the object, emulating a pixel-level dense SSL objective. We count a subcrop hit if it reaches at least $5\%$ object coverage, which we justify as background classes generally have little visual variation and consequently, minimal impact on the pooled representations. We extend this analysis to the BDD100K semantic segmentation dataset, empirically simulating subcrops, and computing subcrop hit and pixel probabilities for varying object sizes. The simulation results, illustrated in Figure 10(b), show a greater relative difference between subcrop hit and pixel probabilities for smaller objects, indicating pseudo-iconic subcrops increase their prevalence in the pooled objective. This likely contributes towards PooDLe’s improved performance on small object classes compared to dense SSL methods. Further details of the analysis can be found in Appendix C.
5 Conclusion
Self-supervised learning on naturalistic videos presents many unsolved challenges, especially due to the presence of high-resolution, multi-object crowded scenes with severe spatial imbalance. Iconic methods rely on single-subject images, and dense methods struggle with the scale imbalance of objects. We propose PooDLe that combines pooled region-invariance learning and dense flow-equivariance learning objectives in a unified framework. PooDLe achieves state-of-the-art performance on downstream semantic segmentation and object detection evaluations compared to prior methods pretrained on the same video datasets, particularly on recognizing small objects. Our study on the effects of crop area, input resolution, and temporal stride also offers key insights on the design choices for video self-supervised learning.
Acknowledgements
We thank Jenny Zhu for her assistance in generating semantic segmentation labels for the WT-Sem dataset. The work is supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) under grant RS-2024-00469482, funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research. AW is supported by the NSERC PGS-D Scholarship. CH is supported by the DoD NDSEG Fellowship. The compute is supported in part through the Microsoft Accelerating Foundation Model Research (AFMR) program, a Google Cloud Platform (GCP) award, and NYU IT’s High Performance Computing resources, services, and staff expertise.
@@ Appendix A Implementation details
As discussed in the pretraining details, we use a Resnet-50 (He et al., 2016a) as our backbone architecture. The projector model is a non-linear, 2-layer MLP (linear for pooled, $1\times 1$ convolutions for dense) that has a $4096$ hidden dimension and projects down to $256$ dimensions. The predictor is the same network with $256-4096-256$ channels. We follow BYOL (Grill et al., 2020) with a momentum starting at $0.996$ and increasing to $1$ throughout training.
@@
Flow is predicted using a supervised off-the-shelf RAFT model or an unsupervised UFlow (Jonschkowski et al., 2020) model that we train ourselves. For unsupervised training, we exactly follow UFlow and train on the KITTI (Geiger et al., 2013) dataset before finetuning on BDD100k (Yu et al., 2020) for 100,000 steps on daytime-only videos. The training and inference resolutions were set to $256\times 512$ to better match the inference setting. KITTI used adjacent frames (10Hz video) while BDD frames were sampled with a temporal stride of 10 (30Hz video).
@@
$K=6$ paired local crops are sampled using the methods described. Cropping is performed using RandomResizedCrop with an output resolution of $192\times 192$. Jitter is 10% of the input image size and a standard aspect ratio range of $[3/4,4/3]$ is used.
@@
We sum our 2 loss functions directly and give them equal weight. The loss computation and warping function were applied to representations after reversing the affine transform and resizing to the input image resolution. This is to take full advantage of high resolution flow like in FlowE (Xiong et al., 2021). We also use flow-based occlusion to prevent misaligning occluded regions without correspondence. We use the same occlusion formulation as DDFlow (Liu et al., 2019) and parameters $\alpha_{1}=0.1,\alpha_{2}=0.5$. We also mask out regions that are not visible after affine transformations for $\mathcal{L}_{\text{dense}}$.
@@
AdamW is used as the optimizer and a weight decay value of $0.01$. A learning rate of $45e-$ 4 is used with 32 GPUs and $4$ image pairs per GPU for a batch size total of $128$. Cosine learning rate decay is used with a schedule for 300 epochs, despite early termination due to compute limitations. LR warmup is used for 2 training epochs. Full float32 precision is used during training.
@@
For all BDD and Cityscapes semantic segmentation and object detection readout tasks, we follow the setup described in FlowE (Xiong et al., 2021) for ResNet-based methods. For ViT-based methods, we adopt those settings, but use AdamW for the optimizer with a learning rate of $53e-$ 5 and weight decay of $0.05$, and a crop size of $512\times 512$ rather than the normal $512\times 1024$ to accommodate the square aspect ratio used in ViT pretraining, following the semantic segmentation linear readout setup described in iBOT (Zhou et al., 2021). In addition, ViT-based methods require sliding window inference in order to achieve performance that is competitive with convolution-based methods.
For ADE20K and WT-Sem linear readout, we simply use the respective BDD linear readout settings for ResNet and ViT methods. For ADE20K and WT-Sem UperNet finetuning, we follow the procedure described in iBOT (Zhou et al., 2021) except we use a batch size of 4 for WT-Sem finetuning.
Ablation data sampling.
For all ablation experiments, we employ repeated sampling like in MAE-st (Feichtenhofer et al., 2022) which samples $R$ frames each time a video is encountered for faster data loading. Therefore, each pass through every video in the dataset counts as $R$ epochs.
@@ Appendix B Compute resources
Ablation-sized experiments were run on 2 or 4 H100/A100 GPUs for a total of 40 epochs, taking 20–40h depending on the configuration.
@@ Appendix C Subcrop analysis
Appendix D Walking Tours Semantic benchmark
(a) Frequency of classes by pixel
(b) Frequency of classes by frame
(c) Unique classes over time
Figure 12: Analysis of WT-Sem in comparison to ADE20K (Zhou et al., 2017) by frequency of each class by pixels occupied (a) or frames present (b) and number of unique classes in each frame (c).
We create the WT-Sem benchmark by sampling a frame every 2 seconds from each of the 10 videos in $\text{WT}_{\text{all}}$ as well as 3 new walkaround videos. The new walkaround videos are filmed in Rome, Torun, and Poznan, sourced from the same YouTube channel as WT (Venkataramanan et al., 2024) under the Creative Commons (CC-BY) license. The Swin-L variant of OpenSeed (Zhang et al., 2023a), pretrained on COCO (Lin et al., 2014) and Objects365 (Shao et al., 2019) and finetuned on ADE20K, is used to generate semantic segmentation masks. We utilize the 25,910 frames sourced from $\text{WT}_{\text{all}}$ as the training set and the 6,170 frames sourced from the 3 new videos as the validation set. Figure 12 shows our analysis of WT-Sem in comparison to ADE20K (Zhou et al., 2017), where we observe that both datasets have long-tailed class distributions and WT-Sem has slightly higher number of unique classes per frame. We also visualize examples from the WT-Sem benchmark in Figure 13.
Figure 13: Visualizations of images and associated semantic segmentation masks from the WT-Sem benchmark.
Appendix E Additional visualizations
We provide additional visualizations of results on our evaluated benchmarks: BDD100K (Yu et al., 2020) semantic segmentation (Figure 14), object detection (Figure 15) and ADE20K semantic segmentation (Figure 16). Once again, we note that PooDLe produces segmentation maps with clearer boundaries while also effectively capturing small objects.
@@ Appendix E Additional visualizations
Appendix F Additional evaluation results
Table 6: Additional BDD semantic segmentation (SemSeg) and object detection (Det) readout evaluations. All settings are conducted with a frozen backbone. BYOL results are taken from FlowE (Xiong et al., 2021) and used DeepLab v3 (Chen et al., 2017) in-place of Upernet (Xiao et al., 2018). *Pretrained on BDD, initialized with supervised ImageNet weights.
@@ Appendix F Additional evaluation results
BYOL (Grill et al., 2020) R50 1000 IN1K 28.3 - 52.4 - 2.8 26.0 39.9 - 60.3 -
DenseCL (Wang et al., 2021b) R50 200 IN1K 21.3 82.7 52.8 91.6 0.3 25.0 27.3 84.0 63.7 93.7
@@
DINO (Caron et al., 2021) ViT-S 800 IN1K 38.5 88.1 52.3 92.0 - - 47.1 90.3 63.6 94.0
iBOT (Zhou et al., 2021) ViT-S 800 IN1K 44.4 89.6 54.2 92.2 - - 52.1 91.5 65.3 94.3
@@
We compare PooDLe against ImageNet-pretrained baselines in Table 6 and observe that PooDLe outperforms most baselines except iBOT and ImageNet supervised ViT. This result is encouraging, as pretraining on naturalistic video is more challenging due to spatial and class imbalance, yet is also a more realistic setting that enables the use of broader sets of usable data. Furthermore, we note that pretraining on class-balanced data such as ImageNet particularly benefits mIoU, which weighs all classes equally despite some classes only appearing in a tiny proportion of pixels in evaluation. Finally, PooDLe pretrained on BDD with weights initialized from the ImageNet supervised checkpoint surpasses all ImageNet-pretrained baselines on linear semantic segmentation.
Appendix G Per-class evaluation results
Table 7: IoU per class on BDD semantic segmentation linear readout. *Pretrained on BDD, initialized with supervised ImageNet weights.
@@ Appendix G Per-class evaluation results
Appendix H Backbone computation cost
We provide a table detailing the number of FLOPs associated with various SSL methods and backbones. We note that our SDM is by far the most efficient upsampling approach for dense representation learning methods.
Architecture Associated Methods GFLOPs
ResNet-50 DINO-R50 43.3
ResNet-50 + SDM PooDLe 60.5
ResNet-50 + FPN decoder PixPro 124.4
ResNet-50 + dilated convolutions FlowE, DINO, DenseCL 200.7
ViT-S/16 DINO, iBOT, DoRA, MAE 82.9
Table 9: Comparison of different backbones, their associated methods, and computational requirements in GFLOPs.
Appendix I Flow visualizations
Figure 17: Comparison of predicted optical flow from RAFT (supervised) and UFlow (unsupervised).
In Figure 17, we compare the predicted flow maps generated from RAFT (Teed & Deng, 2020), an off-the-shelf supervised model, and our own unsupervised UFlow (Jonschkowski et al., 2020) model. The frame pairs are randomly sampled with $\Delta t\in[15,30]$. We do note that self-supervised flow, particularly on BDD100K, may exhibit noisy or splotchy results. This is possibly due to the inconsistent motion and large dark regions that do not offer sufficient photometric supervisory signal. This is in contrast to RAFT (Teed & Deng, 2020) which learns sharp edges like from supervised labels. Nevertheless, we find that this self-supervised flow is sufficient for training PooDLe.